OpenStudy (anonymous):

3a*9a^2+3a*6a+3a*3+3*9a^2+3*6a+3*3 combine like terms and simplify?

OpenStudy (anonymous):

this is just part of my problem

OpenStudy (aravindg):

first can a^2 terms ?

OpenStudy (anonymous):

reiterate plz?

OpenStudy (aravindg):

first can you add a^2 terms ?

OpenStudy (aravindg):

then add a term

OpenStudy (aravindg):

@reason can u do tht ?

OpenStudy (chihiroasleaf):

is this your question ? \[3a \times 9a ^{2} + 3a \times 6a + 3a \times 3 + 3 \times 9a ^{2} + 3 \times 6a + 3 \times 3\]

OpenStudy (anonymous):

27a^3? and to chihiro. yes

OpenStudy (aravindg):

where did a^3 come from ?

OpenStudy (anonymous):

? i'm lost here. i have the rest of the problem

OpenStudy (aravindg):

i asked u to add a^2 terms and a terms

OpenStudy (anonymous):

original problem (3a+3)^3

OpenStudy (chihiroasleaf):

are you asking to expand it?

OpenStudy (anonymous):

then i got (3a+3)(9a^2+6a+3)

OpenStudy (anonymous):

?? then i got what i originally asked and got stuck

OpenStudy (chihiroasleaf):

what's the question ask you to do? expand the expression?